OpenStudy (samigupta8):

Let f is a differentiable function such that f(x)=x^2 + integral of e^-tf(x-t)dt where t goes from 0 to x then f(3) =?

OpenStudy (samigupta8):

@hartnn

OpenStudy (samigupta8):

Actually what i did was first applied Leibnitz's rule and diff f(x). Solving that i got this F'(x)=2x+e^-xf(0) Now f(0)=0 from the equation of f(x) given in the question. So if we just integrate then i got F(x)=x^2 So ans should be 9. Bt it's not ryt.. Where am i wrong?? :(

hartnn (hartnn):

have you tried using the property : \(\large \int_a^b f(z)dz = \int_a^b f(a+b-z)dz\)

OpenStudy (samigupta8):

I didn't.. I just diff the expression n again integrated it!

OpenStudy (samigupta8):

Can u tell is this right . F(0)=0

ganeshie8 (ganeshie8):

thats right \(\int_0^0f(x)\,dx = 0\)

OpenStudy (samigupta8):

So ans is not coming if i do it here.. Like first of all if i diff the expression then i m getting f'(x)=2x+e^-x f(0) That means f'(x)=2x Now just integrate it.. We will get f(x)=x^2 +c Again c=0 for f(0)=0

OpenStudy (samigupta8):

That means f(3)=9.

OpenStudy (samigupta8):

That means f(3)=9.

ganeshie8 (ganeshie8):

your differentiation is wrong

ganeshie8 (ganeshie8):

fundemental theorem of arithmetic 2 gives : \[\dfrac{d}{d\color{red}{x}} \int\limits_a^{\color{red}{x}} g(t)\,dt = g(\color{red}{x})\]

ganeshie8 (ganeshie8):

you have applied it wrong - your integrand contains `x` terms too

OpenStudy (samigupta8):

Okay ! So we can apply this Leibnitz rule only for Integral of g(t)dt containing no other variable..

ganeshie8 (ganeshie8):

\[f(\color{red}{x}) = \color{red}{x}^2+\int\limits_0^{\color{red}{x}} e^{-t}f(\color{red}{x}-t)\,dt\]

ganeshie8 (ganeshie8):

yes, first try and pull the `x` terms out the integrand

OpenStudy (samigupta8):

That can be only done as suggested by hartnn's method!!

ganeshie8 (ganeshie8):

yes, do it

OpenStudy (samigupta8):

Okay!

hartnn (hartnn):

or you can also make a substitution u =x-t

OpenStudy (samigupta8):

I got d ans which is 18..!

OpenStudy (samigupta8):

Thanks !! Both of you

ganeshie8 (ganeshie8):

Awesome!